How many pounds of chocolate worth $1.20 a pound must be mixed with 10 pounds of chocolate worth 90 cents a pound to produce a mixture worth $1.00 a pound?
Answer by mananth(16946)

------------- cost ----------------quantity
chocolate A 0.9 ----------------10
chocolate B 1.2 ----------------x
Total 1 ---------------- 10+x
0.9*10+1.2*x=1(10+x)
9+1.2x=10+x
1.2x-x=10-9
0.2x=1
/0.2
x=5 pounds $1.2 chocolate A